OpenStudy (anonymous):

\[x=3sin\theta\] \[\int x^3 \sqrt (9-x^2)\]

OpenStudy (turingtest):

where are you stuck?

OpenStudy (anonymous):

\[dx=3cos\theta d\theta\] \[9\int \sin^ 3 \theta \cos\theta \sqrt(9-3\sin^ 2 \theta) d\theta\]

OpenStudy (turingtest):

you forgot to square the whole sub\[x=3\sin\theta\implies x^2=9\sin\theta\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

\[9\int \sin^ 3 \theta \cos\theta \sqrt(9-9\sin^ 2 \theta) d\theta\]

OpenStudy (anonymous):

ok...i think I see it now

OpenStudy (turingtest):

further\[x^3=3^3\sin\theta\]so this should be\[\int3^3\sin^3\theta\cos\theta\sqrt{9-9\sin^2\theta}d\theta=27\int\sin^3\theta\cos\theta\sqrt{9-9\sin^2\theta}d\theta\]now we need to get the 9 out of the radical

OpenStudy (anonymous):

\[27\int \sin^ 3 \theta \cos\theta \sqrt(9-9\sin^ 2 \theta) d\theta\]

OpenStudy (turingtest):

\[3^3\int\sin^3\theta\cos\theta\sqrt{9(1-\sin^2\theta)}d\theta=3^3\int\sin^3\theta\cos\theta(3)\sqrt{1-\sin^2\theta}d\theta\]\[=3^4\int\sin^3\theta\cos^2\theta d\theta\]so far so good?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

now it's just u substitution

OpenStudy (turingtest):

I'm no "sir" lol

OpenStudy (anonymous):

no way! LOL! my bad haha

OpenStudy (turingtest):

no it's not because now our cos is to the second power :( so more trouble ahead... not sure what the best move is actually

OpenStudy (turingtest):

reduction formula perhaps

OpenStudy (turingtest):

oh I see what to do :)

OpenStudy (turingtest):

strip out a sin, and convert the sin^2 to 1-cos^2 then u-sub

OpenStudy (anonymous):

sin ^3 = sin * sin ^ 2 = sin * (1 - cos ^ 2)

OpenStudy (anonymous):

yeah thats the best way

OpenStudy (anonymous):

\[=3^4\int\sin^3\theta\cos^2\theta d\theta\] u=cos theta du= sin theta dtheta \[\int (1-cos\theta)cos^ 2 sin\theta d\theta\]

OpenStudy (anonymous):

\[\int (1-u)u^ 2 du\]

OpenStudy (turingtest):

\[3^4\int\sin^3\theta\cos^2\theta d\theta=3^4\int\sin^2\theta\cos^2\theta\cdot\sin\theta d\theta\]\[=3^4\int(1-\cos^2\theta)\cos^2\theta\sin\theta d\theta\]

OpenStudy (turingtest):

somebody dropped a ^2

OpenStudy (anonymous):

there I go again lol

OpenStudy (turingtest):

it happens :)

OpenStudy (anonymous):

\[\int (u^2 - u^4)du\]

OpenStudy (anonymous):

=

OpenStudy (anonymous):

lolz \[27(\frac 1 3 cos\theta- \frac 1 5 cos\theta) \]

OpenStudy (turingtest):

careful with the signs\[u=\cos\theta\implies du=-\sin\theta d\theta\]

OpenStudy (turingtest):

and where did your exponents go ?

OpenStudy (anonymous):

\[27(\frac 1 3 cos^3\theta- \frac 1 5 cos^5\theta)\] I found them...

OpenStudy (anonymous):

wait my negative signs

OpenStudy (turingtest):

yep, now that's your only problem

OpenStudy (anonymous):

\[27(\frac 1 5 cos^5\theta-\frac 1 3 cos^3\theta)\]

OpenStudy (turingtest):

right :) can you put it back in terms of x ?

OpenStudy (turingtest):

wait, it's supposed to be 3^4 on the outside

OpenStudy (turingtest):

oh no wrong again it should be 3^5=243

OpenStudy (anonymous):

where's the 5th one from?

OpenStudy (anonymous):

i think 3^4 is correct

OpenStudy (turingtest):

\[\int x^3\sqrt{9-x^2}dx\]\[x=3\sin\theta\implies du=3\cos\theta d\theta\]\[\int(3\sin\theta)^3\sqrt{9-(3\sin\theta)^2}(3\cos\theta d\theta)\]\[=3^4\int\sin^3\theta\cos\theta\sqrt{9-9\sin^2\theta}d\theta\]\[=3^4\int\sin^3\theta\cos\theta(3)\sqrt{1-\sin^2\theta}d\theta\]\[=3^5\int\sin^3\theta\cos^2\theta d\theta\]

OpenStudy (anonymous):

u=cos x du= sinx then right?

OpenStudy (anonymous):

that would also make it easier to sub the x back in

OpenStudy (anonymous):

i meant u= cos theta

OpenStudy (turingtest):

yeah but you need to use the trig identity first

OpenStudy (anonymous):

well of course \[=3^5\int(1-cos^ 2) \cos^2\theta sin\theta d\theta\]

OpenStudy (anonymous):

I'll be right back

OpenStudy (anonymous):

I gotta go... thank for all of your help again!

OpenStudy (turingtest):

\[\int x^3\sqrt{9-x^2}dx\]\[x=3\sin\theta\implies du=3\cos\theta d\theta\]\[\int(3\sin\theta)^3\sqrt{9-(3\sin\theta)^2}(3\cos\theta d\theta)\]\[=3^4\int\sin^3\theta\cos\theta\sqrt{9-9\sin^2\theta}d\theta\]\[=3^4\int\sin^3\theta\cos\theta(3)\sqrt{1-\sin^2\theta}d\theta\]\[=3^5\int\sin^3\theta\cos^2\theta d\theta\]\[=3^5\int\sin^2\theta\cos^2\theta\sin\theta d\theta\]\[=3^5\int(1-\cos^2\theta)\cos^2\theta\sin\theta d\theta\]\[=3^5\int(\cos^2\theta-\cos^4\theta )\sin\theta d\theta\]\[u=\cos\theta\implies du=-\sin\theta d\theta\implies\sin\theta d\theta=-du\]\[3^5\int(u^4-u^2)du=3^5(\frac15u^5-\frac13u^3)=3^5(\frac15\cos^5\theta-\frac13\cos^3\theta)\]from our initila substitution we get the triangle from which we can find cos in terms of x\[\int x^3\sqrt{9-x^2}dx=3^5\left(\frac15\cdot{(9-x^2)^{5/2}\over3^{5/2}}-\frac13\cdot{(9-x^2)^{3/2}\over3^{3/2}}\right)+C\]which we can simplify a bit